Can someone please help me with this problem?
The larger of two numbers is nineteen more than the smaller. Twice the larger equals ten times the smaller minus ten. What are the numbers?
This is what I have so far and I'm not sure if it is right. Thanks.
Let x= The smaller
Let x + 19 = The larger
2(x=19)= 10(x-10)

Answer by stanbon!(97)

2(x+19)= 10(x-10)
2x+38 = 10x-100
8x = 138
x = 17 1/4
Cheers,
Stan H.